OpenStudy (anonymous):

If the midpoint between (8, y) and (-11, 6) is (-1.5, 5), find the value of y. How do I do this?

OpenStudy (whpalmer4):

The midpoint of a line segment is the average of the coordinates of the endpoints. For example, a line segment from (0,0) to (4,8) has its midpoint at (2,4) because (0+4)/2 = 2 and (0+8)/2 = 4 That means you need to solve \[\frac{y+6}2 = 5\]